OpenStudy (anonymous):

Find the height of the cylinder. V= 271.4 in^3 r= 6 in.

OpenStudy (paki):

simple.... just use the formula... h= v/pie * r^2 h= 271.4/ 3.14 * 6 * 6 h= 2.4 in

OpenStudy (paki):

@Shay17

OpenStudy (anonymous):

Thank you ! :)

OpenStudy (paki):

pleasure :-)